A dad travels 36 miles in 12 hrs. The son travels double in the same amount of time. Where was the son in 9hrs??? I'm guessing the son traveled double so 72 miles in 12 hrs but that about as far as i got. :( thank you much!!!
Answer by lwsshak3(11628)

A dad travels 36 miles in 12 hrs. The son travels double in the same amount of time. Where was the son in 9hrs
..
Dad's rate of speed=36mi/12hrs=3mi/hr
Son's speed as you figured=double that of Dad=2*3=6mi/hr
So in 9 hrs the son should be 9*6=54miles away